OpenStudy (anonymous):

What is i^84? A. -i B. i C. -1 D. 1

OpenStudy (anonymous):

1

OpenStudy (zarkon):

\[i^{84}=i^{4\cdot21}=(i^4)^{21}=1^{21}=1\]

OpenStudy (anonymous):

What about finding the equivalent to \[\sqrt{-121}\] ?

OpenStudy (zarkon):

11*11=121 so \[\sqrt{-121}=11i\]

jhonyy9 (jhonyy9):

- will be (i^2)^42 =(-1)^42 =((-1)^2)^21 =(1)^21 =1

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

the definition of \(i\) is that \[i=\sqrt{-1}\]hence \[i^{84}=(i^{2})^{42}=(-1)^{42}=1\]